### v1

# Adaptive Reinforcement Learning Planning: Harnessing Large Language Models for Complex Information Extraction

Abstract
Existing research on large language models (LLMs) shows that they can solve information extraction tasks through multi-step planning. However, their extraction behavior on complex sentences and tasks is unstable, emerging issues such as false positives and missing elements. We observe that decomposing complex extraction tasks and extracting them step by step can effectively improve LLMs’ performance, and the extraction orders of entities significantly affect the final results of LLMs. This paper proposes a two-stage multi-step method for LLM-based information extraction and adopts the RL framework to execute the multi-step planning. We regard sequential extraction as a Markov decision process, build an LLM-based extraction environment, design a decision module to adaptively provide the optimal order for sequential entity extraction on different sentences, and utilize the DDQN algorithm to train the decision model. We also design the rewards and evaluation metrics suitable for the extraction results of LLMs. We conduct extensive experiments on multiple public datasets to demonstrate the effectiveness of our method in improving the information extraction capabilities of LLMs.
Adaptive Reinforcement Learning Planning: Harnessing Large Language Models for Complex Information Extraction
Zepeng Ding, Ruiyang Ke, Wenhao Huang, Guochao Jiang, Yanda Li, Deqing Yang, Yanghua Xiao, Jiaqing Liang🖂 Shanghai Key Laboratory of Data Science School of Data Science, School of Computer Science, Fudan University {zpding22, ryke23, whhuang21, gcjiang22, ydli22}@m.fudan.edu.cn {yangdeqing, shawyh, liangjiaqing}@fudan.edu.cn
1 Introduction
Current large language models (LLMs) have excellent performance in many fields such as dialogue, reading comprehension, and text generation, greatly expanding the boundaries of traditional NLP tasks and thus demonstrating the capacity to support complex information extraction (IE) tasks. Large models can understand extraction requirements through specific prompts and improve their performance on downstream tasks through in-context learning, multi-step planning Wei et al. (2023), collaborating with small model Ding et al. (2024) and so on. These methods achieve good performance in general IE tasks.
Figure 1: (a) Multi-step extraction reduces errors. (b) Extraction orders have an impact on the outputs.
However, when confronted with complex situations such as very long sentences or paragraphs, multiple related entities in the same sentence, and tokenization confusion caused by overlapping tokens, LLMs still engender issues such as false positives and missing elements Ma et al. (2023); Wang et al. (2024), resulting relatively low F1-score. Fine-tuning LLMs with task data is a common solution to enhance extraction capabilities, but it is time-consuming, has limited generality and its performance highly relies on the quality of the labeled data. Even though LLMs are powerful, they still require multiple steps to achieve acceptable extraction results for complex cases Wei et al. (2023).
Previous studies on small language models (SLMs) show that task decomposition and sequential extraction could reduce errors and achieve better results Xie et al. (2021); Lu et al. (2022). For LLMs, solving the task step by step rather than extracting all the content at once can deconstruct the information of the original sentence, leading to better extraction performance in complicated scenarios. An example is shown in Fig. 1. In our work, we divide the extraction process of LLMs into two distinct tasks: the classification of relations/events and the extraction of entities (arguments). We extract entities in multiple steps, orderly rather than concurrently. Indeed, many studies on LLM-based agents are using large models to solve problems in an orderly step-by-step manner with interaction and feedback, such as ReAct Yao et al. (2022) and AutoCrawler Huang et al. (2024). The effectiveness of these methods heavily depends on the inherent planning ability of LLMs.
Taking complex information extraction as an example, many studies prove that it requires particularly high planning capabilities, which poses a considerable challenge to the aforementioned agents. First, the results of multi-step extraction are sensitive to the order of steps. Previous works use preset prompt templates and present a fixed order planning Wei et al. (2023); Ding et al. (2024). However, motivated by Huang et al. (2023), we find that varying extraction orders significantly influences the output of LLMs. An appropriate order can effectively utilize the logical association between entities and greatly reduce the complexity of the extraction task, thereby driving LLM to obtain better results. Second, the optimal order of entity extraction is not fixed for different sentences, even if the relation/event type is the same. As illustrated in Fig. 1, different extraction orders lead to different results, and even if the relation is the same as that in Fig. 1, the optimal order is distinct for these two sentences. It is difficult to obtain sufficient information to plan the optimal execution order for different sentences Li et al. (2024). Third, the optimal order for models of different capabilities and types is different, making planning and extraction execution entangled and difficult to decompose.
Therefore, we believe that reinforcement learning is essential to generate plans that are acceptable to complex IE cases. In this paper, we establish a framework to effectively guide LLMs for multi-step extraction planning, which can have stable and effective extraction performance in different languages, sentence types, and tasks without fine-tuning. We propose a decision model to assist LLMs in executing entity extraction, adaptively deciding the subsequent action(s) based on the current state. Wed multi-step extractions as a Markov Decision Process (MDP), and design the state based on the input sentence, relation/event type, and extracted contents. We construct the environment with the LLM-based extractor, introduce an indicator as the reward function, and assign reward values based on the semantic correctness and token-level precision of the extraction results. After training through reinforcement learning and subsequently establishing a Deep Q-Network (DQN), the small decision model can collaboratively execute the sequential entity extraction tasks with the LLM-based extractor.
In summary, our main contributions are as follows:
  • First, we design a two-stage multi-step information extraction framework for LLMs, model order selection as MDP, and propose an easy-to-train decision model to guide LLMs in multi-step planning.
  • Second, we adopt RL based framework to train the decision model, which could provide the optimal order for sequential entity extraction. We build the LLM-based extraction environment and design the rewards to reflect the semantic correctness and token-level precision of the results.
  • Third, the experimental results demonstrate the effectiveness of our method. We design an evaluation metric that amalgamates text similarity and exact matching, which is suitable for assessing the extraction capability of LLMs.
2 Related Work
2.1 Research in Information Extraction
In Information Extraction (IE) tasks such as relation extraction and event extraction, research based on traditional models is relatively comprehensive. Pipeline IE methods dissect the extraction process into several sub-tasks, optimizing each individually. For instance, subjects, relations, and objects are sequentially extracted from sentences in pre-specified orders in addressing relation extraction tasks Wei et al. (2020); Xie et al. (2021). When dealing with event extraction, event types, and triggers are initially identified, followed by the extraction of the arguments using sequence tagging models or machine comprehension models Yang et al. (2019). Joint IE methods combine several extraction processes into a single stage. For example, graph-based approaches identify the entity or text span and construct a graph with relations or co-occurrences Wadden et al. (2019); Zhu et al. (2021), while generation-based approaches serialize structured extraction results into sentences or pre-defined templates in a fixed order Hsu et al. (2022); Ye et al. (2021); Lu et al. (2022).
Recent studies on Large Language Models (LLMs) demonstrate their proficient performance across various downstream tasks, even when provided with only a few examples as prompts Agrawal et al. (2022); Jeblick et al. (2023). For IE tasks, some studies indicate that with appropriate prompting, LLMs can achieve comparable performance with the supervised methods in zero-shot or few-shot settings of extraction tasks Wei et al. (2023); Gao et al. (2023); Tang et al. (2023). Ding et al. (2024) integrates LLMs with a small evaluation model for RE tasks, addresses the "low recall" issue, and enhances the extraction performance of LLMs. Furthermore, Chung et al. (2022) shows LLMs can produce outstanding performance through supervision and fine-tuning, and Wadhwa et al. (2023) suggests that LLMs should be a standard benchmark for some extraction tasks.
2.2 Reinforcement Learning in IE Tasks
Previous studies have also employed reinforcement learning (RL) in IE tasks. Narasimhan et al. (2016) enhances information extraction by utilizing RL to integrate external evidence. Takanobu et al. (2019) implements a hierarchical RL framework to enhance the linkage between entity mentions and relation types. Zeng et al. (2018) uses the relation of entity pairs as distant supervision and guides the training of the relation extractor with RL to address large-scale RE tasks. The ordered extraction of entities can be modeled as a sequential decision-making problem, so it can be naturally integrated with RL. For example, Zeng et al. (2019) contemplates the extraction order of relational facts in a sentence and subsequently trains a sequence-to-sequence model using RL, and Huang et al. (2023) proposes an RL-based framework that dynamically generates an optimal extraction order for each sentence to guide the model towards better extraction performance.
3 Methods
3.1 Preliminary and Task Definition
Information extraction (IE) tasks are highly diversified due to their different targets Lu et al. (2022). For this paper, we mainly focus on relation extraction (RE) and event extraction (EE) tasks. RE task aims at extracting triples (subject, predicate, object) (or (s,p,o)) from a given natural language sentence. The subject entity and object entity are linked by a specific predicate (relation). EE task intends to identify the event types from the given sentence and identify the arguments (entities) corresponding to each role. The extracted arguments need to be consistent with the event type and semantically and logically correct.
Figure 2: An example of the multi-step information extraction of EE task.
We uniformly define RE and EE as a two-stage task, as shown in Fig.  2. The goal of the first stage is to classify the relation or event type based on the given sentence, and the second stage is to extract the corresponding entities or arguments. More specifically, the second stage can be formulated as a multi-argument extraction process: given the sentence $s$, the relation/event type $pred$, our goal is to extract all the arguments in $s$ and correctly fill them into the schema $[role_{1},role_{2},...,role_{n}|pred]$. In RE tasks, $n=2$ for subject and object.
Figure 3: An example of modeling the multi-step entity extraction as MDP. In each step, the decision module chooses a role, and the LLM-based environment returns the corresponding argument(s) and reward(s).
3.2 Solution Framework
For different sentences, the order of roles to be extracted should be adaptive rather than fixed. Therefore, we need to consider the previously extracted information instead of directly deciding the extraction order all at once. We train a BERT-based decision model, which determines the role to be extracted next based on the current state (input text and previously extracted information). We regard the order decision as a Markov decision process (MDP) and use reinforcement learning to train the decision model.
Generally, for the multi-step argument extraction task, the LLM-based extractor acts as an environment to provide the decision model with the state after the current extraction execution, and the decision model selects the most appropriate role based on the state to guide LLMs to extract step by step until all roles are extracted.
We use the LLMs in both stages without fine-tuning with labeled data. The output of LLMs is crucial to completing the RL training of the decision model. To make the output of LLMs more stable, and to enable the LLM-based environment to effectively interact with the decision module, we model the reward assignment as a classification task and design a reward model that takes into account both semantic correctness and token-level matching. The reward model and extraction prompts are described in subsequent subsections.
For the classification task, we provide LLMs with input text and a list of candidate relation/event types and also provide several appropriate examples for different datasets in the prompt to guide LLMs to select the type(s) from the candidate list  [^1] LLMs can easily follow instructions to make correct classifications (also verified in Table 4).. For argument extraction in the second stage, each step only requires LLMs to extract arguments (entities) corresponding to a single role. In the $i$-th step, we provide the input text $s$ and the previously extracted role-argument pairs $(r_{j}:a_{j})_{j=1:i-1}$ as input to LLMs, and instruct LLMs to extract the argument $a_{i}$ corresponding to the currently required role $r_{i}$.
3.3 MDP for Argument Extraction
We build the environment with the LLM-based extractor, initialize the state based on the relation/event extracted in the first stage, and determine the roles schema (as the initial action space). In each step, the decision model takes the input sentence and extracted content as input and chooses a previously unselected role as the action. The environment receives the action from the decision model, constructs the extractor’s input, and performs extraction. After collecting the extraction results, the environment would assign the reward based on the ground truth of the labeled sample, separate the different arguments of the current role, build the successor state(s), and transit to a new state. When all roles are selected (the action space is empty), the decision model returns a termination signal, and the environment converts the extracted content into structural output. Fig. 3 shows the entire extraction process of an instance.
State
The state includes the original sentence $s$, relation/event $p$, extracted content $C$, and roles schema $\mathcal{M}$. We denote the state of step $i$ as $S_{i}$, as follows:
$S_{i}=(\mathcal{M},C_{i},p,s)$ (1)
Where the extracted content consists of previously extracted role-argument pairs: $C_{i}=(r_{j}:a_{j})_{j=1:i-1}$. In each step, the environment constructs the input of LLM-based extractor according to the current state $S_{i}$ and the role $r_{i}$ selected by the decision model and transits to a new state.
Action
For a given sentence and corresponding relation/event type, the action of the decision model is the next role to extract. The initial action space $A_{0}$ is the roles schema $\mathcal{M}$ in the state. In a complete argument extraction process, the extraction order is a permutation of roles in $\mathcal{M}$, and each role should be extracted only once. Therefore, the action space of the decision model is reduced at each step. After selecting action (role) $a_{i}=r_{i}$ at the $i$-th step, $a_{i}$ will be removed from the action space $A_{i}$ to derive $A_{i+1}$. The update of the action space can be expressed by the following formula:
$A_{i}=\begin{cases}\mathcal{M},&i=0\\ A_{i-1}-\{a_{i-1}\},&0<i<|\mathcal{M}|\end{cases}$ (2)
Reward
The design of rewards is crucial in RL training and is directly related to the loss and the update of model weights. In our framework, the reward value is used to measure the quality of the extraction results and guide the decision model to make better plans for each episode. However, the output of a large model is usually uncertain and imprecise. On one hand, applying the exact match with the ground-truth will misjudge many semantically correct answers as wrong, reducing the effectiveness of the reward value. On the other hand, IE tasks require accurate results in some specific areas and scenarios. Therefore, we hope to assign high reward values to outputs that are both semantically correct and highly matched with the ground-truth, and use low reward values to guide LLMs to avoid outputs that are semantically incorrect or mismatched with the ground-truth. Details of our reward function and reward model are as follows.
3.3.1 Reward Function and Reward Model
Based on the above considerations, we set the reward function to an indicator function, that is, to assign a reward value of 1 to acceptable extraction results [^2] To match the magnitude of the $Q$ value generated by the decision model, in practical we use 0 or 10 as the reward., and to assign a reward value of 0 to semantically incorrect or mismatched answers. We directly use an LLM to evaluate the semantic correctness, automatically implement this concise but effective reward function, and use token-level similarity between the extracted results and the ground-truth as a threshold to help determine whether they match.
Note that our reward function has only two discrete values (which can be regarded as two labels), rather than a continuous function. On one hand, it is intuitive and reasonable to classify the extraction results into "acceptable" and "need to be rejected", but it is difficult to find a reliable and stable method to further sort the acceptable results and assign different rewards. On the other hand, using an LLM as our reward model is convenient and effective. It only needs to judge whether the extraction results are acceptable and output 0 or 1.
During the RL training process, we use an LLM with a larger scale and stronger capabilities than the extraction model as the reward model to ensure accurate and effective reward assignment. This process also performs implicit knowledge distillation through the rewards. We instruct it to consider whether the semantics of the extracted arguments are consistent with the ground-truth and the token-level accuracy. See Appendix C for the prompt template.
We use the word-level textF1 score Huang et al. (2023) defined as follows to measure the token-level similarity between ground-truth and extracted arguments and use it as a threshold to remove "mismatched" extraction results. This step can be regarded as a pruning process, which reduces the successor state space and enhances the effectiveness when performing inference.
$Set_{a}=$ $Set_{p}\cap Set_{g}$ (3)
$Text\text{P}=\frac{|Set_{a}|}{|Set_{p}|},$ $\quad Text\text{R}=\frac{|Set_{a}|}{|Set_{g}|}$
[…]
3.4 Training of Decision Model
We use a BERT-based decision model to evaluate the value of (State, Action) pair to provide the next action for the environment. Specifically, we use BERT Devlin et al. (2019) for English datasets and RoBERTa Liu et al. (2019) for Chinese datasets. We concatenate the input state and action into strings to form a sequence $x_{t}$ = [[CLS],$\mathbf{a_{t}}$,[SEP],$\mathbf{S_{t}}$,[SEP]].
The Bert encoder converts these tokens into hidden vectors, and we select $h_{0}$ (the encoding vector of the token [CLS]) as the representative of the (State, Action) pair. The final evaluation of $Q(S_{t},a_{t})$ is defined as:
$\hat{Q}=\mathbf{Wh_{0}}+\mathbf{b}$ (4)
where $\mathbf{W}$ and $\mathbf{b}$ are trainable parameters.
Input: $\mathcal{D}:$initialized replay memory; $\theta$: network parameters; $\theta^{\prime}$: target-net parameters; $N_{b}$:training batch size
Input: hyper-parameters $E,\epsilon,\gamma,k$
1for epoch=1,…,$E$ do
2       Sample instances from labeled train set.
3       Initialize the state $S_{0}=(\mathcal{M}_{0},C_{0},p,s)$.
4       for $|\mathcal{M}|$ do
5             $p=$ Random$(0,1)$
6             if p<1-$\epsilon$ then
7                   $a_{t}=\arg\max_{a}Q(S_{t},a;\theta)$
8            else
9                   $a_{t}=$ Random-sample$(\mathcal{A}_{t})$
10             end if
11            $S_{t+1},r_{t}=$ Transit$(s_{t},a_{t})$
12             Store $(S_{t},a_{t},r_{t},S_{t+1})$ in $\mathcal{D}$
13             Sample random mini-batch of $N_{b}$ transitions $(S_{j},a_{j},r_{j},S_{j+1})$ from $\mathcal{D}$, and do follows:
14             if $S_{j+1}$ is terminal state then
15                   $y_{j}=r_{j}$
16            else
17                   $y_{j}=r_{j}+\gamma\max_{a}Q(S_{j+1},a;\theta^{\prime})$
18             end if
19            Update parameter $\theta$ on the loss $\mathcal{L}(\theta)=(y_{j}-Q(S_{t},a_{t};\theta))^{2}$
20             Replace target-net parameters $\theta^{\prime}$ = $\theta$ every $k$ steps
21       end for
22
23 end for
Algorithm 1 Training of decision model: DDQN algorithm with experience replay
We apply Deep Q-Learning to train the model and build a Deep Q-Network (DQN). Unlike the traditional DQN Mnih et al. (2013), in our tasks, a role may correspond to multiple arguments, and the extractor could extract multiple answers (arguments) based on the current state, or may not return any results (if there is no argument that matches the extracted content and the role). We design the following learning object:
$Q(S,a)=r_{(S,a)}+\gamma\frac{1}{|\mathcal{S}_{(S,a)}|}\cdot\sum_{S^{\prime}\in\mathcal{S}_{(S,a)}}\max_{a^{\prime}\in\mathcal{A}}Q(S^{\prime},a^{\prime})$ (5)
where $r_{(S,a)}$ is the reward, $\mathcal{S}_{(S,a)}$ is the set of the successor states derived from the state $S$ with action $a$, and $\gamma$ is the discount factor.
We adopt the DDQN algorithm Van Hasselt et al. (2016), utilize $\epsilon-$greedy exploration and the Experience Replay Mnih et al. (2013) for RL training, as illustrated in Algorithm 1. The loss function is defined as the expected value of the mean squared Temporal Difference (TD) error:
$\mathcal{L}=\mathbb{E}(\hat{Q}(S,a)-Q(S,a))^{2}$ (6)
Furthermore, for different datasets of the same language and task type (such as different English RE datasets), we only need to mix the training data to train one decision model, which simplifies the training process of the decision model and improves the generalization of our method.
3.5 LLM for Relation/Event Classification
We directly apply the LLMs as a classifier, which is more straightforward and simpler than training a BERT-based classifier and does not necessitate the provision of additional information such as trigger words. All potential relation/event types within the dataset are compiled into a list and supplied as input, in conjunction with the input text.
In addition, in some datasets, the relation/event types of the originally labeled data are not natural language expressions. Consequently, it is challenging for LLMs to comprehend the semantics of these relation/event words without undergoing fine-tuning training. To address this, we undertake a natural language conversion for these labels, such as converting "people-deceased_person-place_of_death" in NYT10 dataset to "dead in". Please refer to Appendix B for details, and Appendix C for prompt template.
3.6 LLM-Based Argument Extractor
We adopt an LLM-based extractor to solve the argument extraction task with different extraction orders. For EE tasks, the roles are determined by the event type, and the event type has a corresponding roles schema. The extractor extracts the arguments for the selected role. For RE tasks, the intuitive approach is to set two roles for all relation types: subject and object. However, unlike roles in EE, the words "subject" and "object" do not contain any information about the entity type, which sometimes leads to ambiguity. For example, the subject-object of the relation "affiliated" is person-organization in some datasets, while in other datasets it is department-organization. Therefore, we design the roles in RE tasks as "subject/object: entity type". The entity type is determined by the semantics of the selected relation type in the specific dataset. See Appendix B for examples.
The input of the arguments extractor is a dictionary, including the original sentence, the current relation/event type, the extracted content (role-argument pairs), and the role (argument type) to extract. At each step of argument extraction, the extractor returns the arguments corresponding to the specified role in a string, separated by commas. The prompt template is provided in Appendix C.
4 Experiments
4.1 Datasets
We evaluate our methods on several public and accessible complicated information extraction datasets, including the relation extraction datasets NYT, NYT11-HRL, Wiki80, SKE21, HacRED and DuIE, and the event extraction dataset DuEE and ACE05. These datasets are challenging for LLMs and widely used for evaluation of existing extraction methods. A brief introduction to these datasets is provided in Appendix A.
4.2 Comparing Methods and Metrics
We select some recently popular and high-performance LLMs as extractors in our experiments. We choose the open-source model mistral7B-instruct-v0.3333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Qwen1.5 Bai et al. (2023), and the famous GPT3.5-turbo Brown et al. (2020); Ouyang et al. (2022) for both English and Chinese datasets. We apply two previous multi-stage extraction methods to our testset, ChatIE Wei et al. (2023) and LLM(without fine-tuning)+Filter Ding et al. (2024) and compare our method with them. We also compare our method with random/fixed selection of actions, and specifically calculate the metrics in cases involving multiple triples or roles.
We design a relaxed evaluation method based on the exact match. We consider the correctness of the arguments at the word level and calculate the similarity between the extraction result and ground truth for each argument (according to Eq. 3). In the level of the entire triple or event, we believe the extraction result matches the ground truth only if the relation/event type is classified correctly and the word-level similarity is greater than the threshold for all arguments. We calculate and report precision (Prec.), recall (Reca.), and F1 scores for all the experiments according to the method above. We repeat the experiment 3 times and averaged the metrics as final results. See Appendix D for details.
4.3 Effectiveness of LLM Classification
We first assess the efficacy of LLM-based relation/event classification on the SKE21, DuEE and ACE05 datasets. Table 4 in Appendix E shows that LLMs classify the relation/event accurately and achieve stable and accurate results, setting the stage for the argument extraction.
HacRED SKE21 DuIE DuEE
Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo – ChatIE 81.61 48.51 60.85 71.67 43.31 53.99 70.13 66.13 68.07 57.05 78.13 65.95
– Filter 22.34 25.89 23.98 43.08 51.27 46.82 58.38 72.56 64.70 - - -
– ordered -RL (ours) 32.93 32.72 32.82 45.84 53.64 49.43 63.53 66.99 65.21 58.95 79.58 67.73
-46.06% -8.45% -4.20% +2.70%
Qwen1.5-14B – ChatIE 55.99 19.85 29.31 48.94 35.25 40.99 47.40 39.14 42.87 49.19 38.06 42.91
– Filter 22.34 25.89 23.98 39.86 57.11 46.95 26.42 48.08 34.10 - - -
– ordered -random 31.60 34.82 33.13 67.10 70.51 68.76 31.13 46.99 37.45 53.39 60.14 56.56
– ordered -fixed 29.82 29.07 29.44 60.68 63.31 61.97 28.47 40.75 33.52 50.38 55.83 52.96
– ordered -RL (ours) 36.02 39.98 37.89 76.01 79.75 77.83 40.63 57.10 47.47 53.63 62.64 57.78
+14.37% +13.19% +9.10% +2.11%
Mistral-7B-v0.3 – ChatIE 34.56 7.09 11.76 31.50 8.73 13.67 42.84 31.18 36.09 37.73 11.94 18.14
– Filter 6.03 14.37 8.49 28.81 38.60 32.99 36.53 37.27 36.90 - - -
– ordered -RL (ours) 10.40 12.07 11.17 27.04 32.29 29.43 40.05 36.34 38.10 54.21 69.72 61.00
-5.02% -10.79% +3.25% +236.27%
Table 1: The main evaluation results of different methods on Chinese IE datasets. Better precision, recall and F1 score are marked bold. We also report the percentage improvement on F1 score.
NYT10 NYT11-HRL Wiki80 ACE05
Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo – ChatIE 46.27 42.70 44.41 32.12 42.97 36.76 32.14 29.57 30.80 57.52 22.18 32.02
– Filter 39.12 44.06 41.44 33.06 47.19 38.88 17.08 18.15 17.60 - - -
– ordered -RL (ours) 52.31 68.84 59.44 32.98 49.19 39.49 31.90 48.29 38.42 25.57 65.19 36.73
+33.84% +1.57% +21.94% +14.71%
Qwen1.5-14B – ChatIE 39.66 15.66 22.45 27.48 19.46 22.78 16.67 18.29 17.44 41.38 12.29 18.95
– Filter 12.36 19.22 15.04 25.82 29.97 27.74 6.64 9.50 7.82 - - -
– ordered -RL (ours) 62.62 45.54 52.73 30.68 35.41 32.87 31.18 36.57 33.66 39.58 77.82 52.47
+134.88% +18.49% +93.00% +176.89%
Mistral-7B-v0.3 – ChatIE 25.61 11.84 16.19 19.38 13.51 15.92 6.18 9.14 7.37 51.16 47.51 49.27
– Filter 38.50 59.91 46.88 16.67 19.41 17.94 8.34 9.92 9.06 - - -
– ordered -random 49.73 83.22 62.26 23.57 54.18 32.85 18.18 35.43 24.03 28.03 81.23 41.68
– ordered -fixed 47.87 80.82 60.13 23.26 59.73 33.48 16.01 32.00 21.34 30.39 89.42 45.37
– ordered -RL (ours) 50.47 85.02 63.34 23.70 59.19 33.85 20.43 40.29 27.12 39.94 94.88 56.22
+1.73% +1.11% +12.86% +14.11%
Table 2: The main evaluation results of different methods on English IE datasets. Better precision, recall and F1 score are marked bold. We also report the percentage improvement on F1 score.
4.4 Main Results
Table 1 and  2 show the extraction results on the Chinese and English datasets respectively. ChatIE Wei et al. (2023) uses a multi-step instruction to guide chatgpt to extract, we also apply this instruction template to the open source model and report the results, as shown in the first row of each block. Ding et al. (2024) adds the filtered results of small models to the prompt of LLMs, and we report the performance of this method on our test set, as shown in the second row of each block. All the LLMs in the table have not been fine-tuned on the labeled train set.
The results indicate that, in most cases, our method has better extraction capabilities than the other two prompt-based fixed-ordered planning methods. Previous methods usually only get good results on specific models or datasets, and lack generalization ability. For example, the results of ChatIE on open-source models are far worse than the results on GPT3.5-turbo, and in most cases when ChatIE achieves the best results, our method also gets high recall and F1 score.
In general, for various types of LLM extractors and IE datasets, our method can achieve the highest precision, recall, and F1 score in most cases. Our RL-based multi-step planning method is effective and generalizable.
4.5 Results of Different Step Orders
In addition, we also compared the extraction results of different order selection methods for entity extraction, using Qwen for Chinese data and Mistral for English data, as shown in Table 1 and Table 2. In most cases, RL-based extraction order can achieve the best results, and the metrics of fixed extraction order are lower than random order. These results imply that the optimal order varies across distinct instances, and our method could adaptively make the best plan.
4.6 Complicated Extraction Settings
To further demonstrate the stability and effectiveness of our method and reflect the advantages of dynamic RL-based order selection, we choose more complex situations for experiments, which contain multiple triples or roles. We select sentences with more than 4 triples from the SKE21 test set, and events with more than 5 roles from the DuEE test set to conduct experiments.
The results of Table 5 in Appendix E show that when the IE task becomes complicated, the extraction performance of ChatIE is far inferior to our method. In addition, compared with random or fixed extraction order, our method achieves significant improvements in the metrics. This also suggests that when LLMs undertake extraction tasks in complex scenarios, an appropriate order can further augment the model’s extraction ability.
5 Conclusion
In this paper, we propose a two-stage multi-step method for information extraction tasks of LLMs. We use LLM extractors as the environment, design a reward module that takes into account both semantic correctness and token-level matching, and employ a reinforcement learning framework to train the decision model, which can adaptively select the optimal extraction order.
Experimental results show that our method performs better than fixed-ordered planning in most cases and can be applied to various LLMs, consistently enhancing their IE capabilities. By comparing different order selection methods, it is also verified that our RL-based framework can adaptively perform multi-step extraction planning, bringing stable effect improvement in both general and complex situations.
Limitations
Despite the effectiveness of our model, there are still some limitations of our method.
First, our reward model is concise and effective. In most cases, it is not difficult for a large model to make binary reward assignments based on the semantics of extraction results and the ground-truth. However, it only considers the correctness of each step and lacks the evaluation of the complete extraction results of the entire sentence. Subsequent research can further explore the design of rewards of LLM-based environments.
Second, the decision module of sequential argument extraction in our framework is still based on small models. Furthermore, the training of decision model requires extensive interactions with LLMs, which is relatively time-consuming. In future research, we could contemplate merging similar datasets or texts to train the decision model, or further explore approaches for LLMs to directly undertake the order-decision tasks.
Ethics Statement
We hereby declare that all authors of this article are aware of and adhere to the provided ACL Code of Ethics and honor the code of conduct.
Use of Human Annotations
Human annotations are only used in methodological research at the beginning of the work, to assist in analyzing the feasibility of the proposed solution. Annotators consented to the use of data for research purposes. We ensure that the privacy of all annotators is protected throughout the annotation process, and all of them are adequately paid according to local standards. Human annotations are not applied during the evaluation of our method.
Risks
In this paper, all datasets are obtained from public sources. The datasets adopted have been anonymized and do not contain offensive information. However, we cannot guarantee that the datasets do not contain socially harmful or toxic language.
Appendix A Dataset Introduction
NYT series
NYT is based on the articles in New York Times. There are many derived datasets with better labeling. NYT10 Riedel et al. (2010) and NYT11 Hoffmann et al. (2011) label the complete entities. Moreover, NYT10-HRL and NYT11-HRL Takanobu et al. (2019) are better versions that are processed by optimizing the relation labels.
Wiki80
Wiki80 Han et al. (2019) is a sentence-level dataset based on the few-shot dataset FewRel Han et al. (2018), which contains 80 types of relations. The number of each relation is 700, with a total of 56,000 samples.
HacRED
HacRED Cheng et al. (2021) [^4] https://github.com/qiaojiim/hacred is a novel challenging extraction dataset. It analyzes the performance gap between popular datasets and practical applications, and carefully selects and designs more hard cases. HacRED consists of 65,225 relational facts annotated from 9,231 wiki documents with sufficient and diverse hard cases, which poses a very high challenge to many current complex extraction methods.
SKE21
SKE19555http://ai.baidu.com/broad/download?dataset=sked is published by Baidu, and is currently the largest dataset available for complex relational triple extraction. Since its testing set is unpublished, and there are some errors in the validation set, a version named SKE21 is published by Xie et al. (2021). The testing set of SKE21 is carefully manually relabeled and contains 1,150 sentences and 2,765 annotated triples.
DuIE
DuIE Li et al. (2019) is a Chinese information extraction dataset. It contains more than 210, 000 sentences and 48 pre-defined schema gathered from Baidu Encyclopedia, Baidu Tieba and Baidu Information Stream. Compared to the previous version, it contains 5 Multiple-O-values schema (Schema with multiple object slots and values), which greatly increase the difficulty of the task.
DuEE
DuEE Li et al. (2020) is a Chinese document-level event extraction dataset. It contains 11, 224 documents categorized into 65 event types, along with 41, 520 event arguments mapped to 121 argument roles, which is the largest Chinese EE dataset.
ACE05
ACE 2005666https://catalog.ldc.upenn.edu/LDC2006T06 Multilingual Training Corpus was developed by the Linguistic Data Consortium (LDC) and contains approximately 1,800 files of mixed genre text in English, Arabic, and Chinese annotated for entities, relations, and events. In this paper we use the English events corpus, it provides event annotations in document and sentence levels from a variety of domains such as newswires and online forums.
Appendix B The Inputs of LLMs
In Sections 3.5 and  3.6, we introduce the inputs for the LLM-based classifier and extractor respectively. We describe the natural language conversion of relation/event types, and the design of roles in RE tasks when extracting arguments. Here we provide some illustrative intuitive examples based on the NYT11 dataset.
B.1 Natural Language Conversion of Types
We convert the original relation words in the dataset into natural language expressions that are conducive to the comprehension of LLMs. For example,
    /location/location/contains -> location contains
    /business/person/company -> work at
    /people/person/place-of-birth -> born in
    /business/company/founders -> founder
B.2 Roles in RE Tasks
We design the roles in RE tasks as "subject/object: entity type" to reduce the ambiguity of relation types and promote the argument extraction of LLMs. For example, in NYT11 datasets:
    the roles of relation "location contains" are: "subject: place" and "object: place"
    the roles of relation "work at" are: "subject: person" and "object: company"
    the roles of relation "born in" are: "subject: person" and "object: place"
    the roles of relation "founder" are: "subject: company" and "object: person"
Appendix C Prompt Templates
Prompt template for classification:
Given the origin sentence, and a relation/event type list $L$, Please select the relation/event contained in this sentence from the list $L$ If more than one relation/event in $L$ is included in this sentence, please separate them with commas(’,’) and output.
Here are some examples:
Input: $s_{1},L_{1}$
Output: $o_{1}$
Now given the following sentence and relation/event list, please select all relations or events contained in the sentence, and output them.
Note: Consider only the types provided in the list.
Prompt template for entity extraction:
Given the text $s$ and the interested relation/event $p$, we may also give the extracted content $C$. Please identify the argument of the role $r$ according to the given content.
Here are some examples:
Input: {$s_{1},p_{1},C_{1}$, "the role I want": $r_{1}$}
Output: $a_{1}$
Now given the following input, please complete the extracting task.
Note: The output content should only contains the arguments of the role I want, and needs to match the relation/event type and extracted content provided in the input. If there are multiple corresponding arguments, please separate them with commas(’,’) and output.
Prompt template for reward assignment:
Now there is a model extraction competition. Please judge whether the answer extracted by the model is correct based on the ground-truth and give a score. If the extraction is correct, please give a score of 1, otherwise 0.
Here are some examples:
Extracted: USA; Ground-truth: the United States; Output: 1
Extracted: Beijing City; Ground-truth: Nanjing City; Output: 0
You only need to output the score value 1 or 0, without any redundant output.
Note: The extraction result does not have to be exactly the same as the ground-truth. It can be judged as correct if the meaning is consistent with the ground-truth and does not cause confusion.
Please give a score to the following extracted answer.
Appendix D Experiment Details
Our experiments are conducted on three A800 GPUs. All deep models, including the extraction model and decision model, are implemented using the PyTorch framework. For all parts involving the LLMs, we provide 3 examples in the prompt, directly using LLMs to extract without fine-tuning with labeled data. The time out for all LLMs is set to 6s. As for the RL training of the decision module, we set the buffer size to 5000 and the target network update step to 20. For the reward model, we choose Qwen1.5-72B for both Chinese and English datasets.
Hyper-parameters LLMs Decision Model
Batch size - 32
Max length 2048 512
Learning rate - 1e-4
Temperature 0 -
Top p 1 -
Table 3: Hyper-parameters for LLMs and decision model.
To reduce time consumption, for each dataset, we randomly select 2000 items from the training set for the training of the decision model. We train 10 epochs for all datasets. Additionally, the exploration parameter $\epsilon$ is initialized at 0.9 and the discount factor $\gamma$ is set to 0.5. The exploration rate $\epsilon$ becomes 0.9 times its own in every 100 steps until it reaches 0.05.
We employed AdamW optimizer as the optimizer, with using linear scheduling with warming up proportion 10%. For testing, we randomly select 800 items from the test set for each dataset, and the test data is the same for all methods and LLMs. More details are listed in Table 3. Our codes and prompts can be found at https://anonymous.4open.science/r/Ordered-LLM-IE-E89D.
SKE21 DuEE ACE05
Prec. Reca. F1 Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo 88.68 91.12 89.88 87.92 93.57 90.66 91.93 92.50 92.21
Qwen1.5-14B 91.49 93.93 92.69 87.69 88.58 88.13 85.25 89.00 87.08
mistral7B-instruct-v0.3 79.05 88.42 83.47 81.40 87.43 84.31 88.91 90.16 89.53
Table 4: The results of relation/event classification of different LLMs on SKE21, DuEE and ACE05.
SKE21 DuEE
Prec. Reca. F1 Prec. Reca. F1
GPT3.5-turbo + ChatIE 17.37 28.56 21.60 42.31 65.38 51.37
+ ordered -random 19.07 33.87 24.40 55.48 80.57 65.71
+ ordered -fixed 6.62 11.75 8.47 48.16 67.14 56.09
+ ordered -RL (ours) 24.58 40.09 30.47 59.39 86.62 70.47
+24.88% +7.24%
Qwen1.5-14B + ChatIE 25.31 14.06 18.07 48.72 23.75 31.93
+ ordered -random 55.44 63.36 59.14 60.23 67.52 63.66
+ ordered -fixed 49.79 55.76 52.61 57.27 61.46 59.29
+ ordered -RL (ours) 63.69 72.35 67.75 65.19 79.94 71.82
+14.56% +11.36%
Table 5: The extraction results of complicated situations on SKE21 and DuEE. We select sentences with more than 4 triples from the SKE21, and events with more than 5 roles from the DuEE.
Appendix E Additional Results
Here we present some experimental results as a supplement to Section 4. Table 4 shows the results of relation/event classification of different LLMs, and table 5 presents the results of complicated scenarios on SKE21 and DuEE.